OpenStudy (anonymous):

simplify the product using foil (6x-5)(2x-3)

OpenStudy (anonymous):

@praxer

OpenStudy (anonymous):

12x^2+2x+15

OpenStudy (jadeishere):

6x * 2x = 12x^2 6x * -3 = -18x -5 * 2x = -10x -5 * - 3 = 15 12x^2 - 18x - 10x + 15 12x^2 - 28x + 15 Either i went way wrong or...

OpenStudy (anonymous):

i need the work

OpenStudy (jadeishere):

what do you mean?

OpenStudy (anonymous):

thank u

OpenStudy (jadeishere):

lol okay